Jane is two times as old a jim.
Jim is n years old now.
How old will Jane be in three years?
There is no numerical answer.
I need an expression to show how to solve it.

Answer by stanbon(75887)

Jane is two times as old a jim.
Jim is n years old now.
How old will Jane be in three years?
-----
Now Equations:
a = 2j
j = n yrs old
-------
In 3 years Equation:
(a+3) = 2(j+3)
a + 3 = 2j + 6
a + 3 = 2n + 6
-------
a = 2n+3 (Jane's age NOW)
===============================
Cheers,
Stan H.
=================